The hypotenuse of a given right triangle is 6 cm longer than the
shorter leg. The length of the shorter leg is 2 cm less than that of the longer
leg. Find the length of the three sides

Answer by richwmiller(17219)

I guess I can't help you. I fail to see the finance involved.
c=6+a
a=b-2
a^2+b^2=c^2
a = 4*(1+sqrt(3)), b = 2*(3+2sqrt(3)), c = 2*(5+2sqrt(3))